IzukuMidoriyia:

Determine whether m = 27 is a solution to 20 = m + 3 Yes No

supie:

So basically all you do is plug in the value of `m` in \(20 = m + 3\) so \(20 = 27 + 3\) So does 27+3=20?

IzukuMidoriyia:

no

Noonaaaaaa:

No m equals 17

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @IzukuMidoriyia no \(\color{#0cbb34}{\text{End of Quote}}\) Exactly so `no` is your answer.

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Noonaaaaaa No m equals 17 \(\color{#0cbb34}{\text{End of Quote}}\) Yes correct but try not to give direct answers please...

IzukuMidoriyia:

thx for the help

supie:

np

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie So basically all you do is plug in the value of `m` in \(20 = m + 3\) so \(20 = 27 + 3\) So does 27+3=20? \(\color{#0cbb34}{\text{End of Quote}}\) good job

supie:

ty so much

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie ty so much \(\color{#0cbb34}{\text{End of Quote}}\) you r good like a math teacher - congrats !

supie:

ty

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie ty \(\color{#0cbb34}{\text{End of Quote}}\) np yw